I toss an unfair coin 12 times. This coin is 65% likely to show up heads. Since the binomial table only gives up to 50%, I will have to look at probability for tails. Which column will I be looking in (ie. .05, .30, .50...)?
Answer #1

P(getting tail) = 1-P(getting head) = 1-0.65 = 0.35